OpenStudy (unklerhaukus):

Simplify : \[=\frac{1}{4(b-a)^2}\left(b^3-3a^2b+6a^3\right)\] NB: has a nice solution

OpenStudy (unklerhaukus):

should i try; partial fractions, polynomial long division, some other complicated method, .../?

OpenStudy (unklerhaukus):

i am not confident in these methods

OpenStudy (kinggeorge):

First guess is to use polynomial long division.

OpenStudy (unklerhaukus):

bother, polynomial long division is not my strong suite.

OpenStudy (kinggeorge):

If I use long division I get\[{1 \over 4}\cdot\left({{4a^3 \over {(b-a)}^2}+2a+b}\right)\]Which is a little bit nicer.

OpenStudy (unklerhaukus):

the solution in the back of my book says \[=\frac{2a+b}{4}\]

OpenStudy (kinggeorge):

Well, \[{a^3 \over {(b-a)}^2} \neq 0\]so it looks like they left out a term to me.

OpenStudy (unklerhaukus):

yeah the solution could be incorrect,

OpenStudy (unklerhaukus):

\[=\frac{1}{4(b-a)^2}\left(b^3-3a^2b+6a^3\right)\] \[=\frac{1}{4} \times (b-a)^2 \overline {\bigg)b^3-3a^2b+6a^3}\] \[=\cdots \]

OpenStudy (unklerhaukus):

can you tell me the steps i always forget

OpenStudy (kinggeorge):

First FOIL \((b-a)^2=b^2-2ab+a^2\). Then let's divide with respect to \(b\) first. so we need to find what we multiply \(b^2\) by to get \(b^3\). This is \(b\). So that's the first thing we put on top of the division symbol thing.

OpenStudy (unklerhaukus):

\[\quad\qquad\qquad\qquad b\]\[=\frac{1}{4} \times (b-a)^2 \overline {\bigg)b^3-3a^2b+6a^3}\]\[\quad\qquad\qquad\qquad b^3-2b^2a+a^2b\]

OpenStudy (unklerhaukus):

\[\quad\qquad\qquad\qquad b\]\[=\frac{1}{4} \times (b-a)^2 \overline {\bigg)b^3-3a^2b+6a^3}\]\[\quad\qquad\qquad\qquad b^3-2b^2a+a^2\]\[\quad\qquad\qquad\qquad \quad-3a^2b+2b^2a+5a^2\]

OpenStudy (kinggeorge):

So now you have (with some steps not included) \[={1 \over 4}\left(b+ {2ab^2-4a^2b+6a^3 \over b^2-2ab+a^2}\right)\]Since \[b^3-3a^2b+6a^3 - (b^3-2b^2a+a^2b)=2ab^2-4a^2b+6a^3\]Remember you're subtracting an \(a^2b\) not just an \(a^2\)

OpenStudy (unklerhaukus):

\[\quad\qquad\qquad\qquad b-\frac{3}{2} a \]\[=\frac{1}{4} \times (b-a)^2 \overline {\bigg)b^3-3a^2b+6a^3}\]\[\quad\qquad\qquad\qquad b^3-2b^2a+a^2\]\[\quad\qquad\qquad\qquad \quad-3a^2b+2b^2a+5a^2\]\[\quad\qquad\qquad\qquad \quad \frac{3}{2}ab^2-3a^2b+\frac{3}{2}a^3\]

OpenStudy (kinggeorge):

Now that we have this new equation, do long division one more time with respect to a this time. If I've done everything correct, it should be what we're looking for.

OpenStudy (kinggeorge):